OpenStudy (anonymous):

A line passes through the point (-4,1) and has a slope of 2/3. If you use this slope to generate another point on the line, what are the coordinates of that point? (-1,3) (-2,4) (-12,2) (-8,3)

OpenStudy (anonymous):

the coordinates that satisfy the equation \[slope = y2-y1 / x2-x1\]

OpenStudy (anonymous):

\[2/3 = y - 1 / x + 4\] Try all coordinates in this equation